4.a) Let R[x] deg≤2 be the vector space of all polynomial functions in a single variable x with real coefficient and degree at most 2

Let R[x] deg≤2"\to"R[x] deg≤2 be defined by T(p(x)) =p(x-1)

Let U = <1,x,x2> be the usual basis for R[x] deg≤2 and B = <1+x+x2,2x+x2,x+x2> be another basis for R[x] deg≤2

Let, for any polynomial p(x) "\isin" R[x] deg≤2, coordB(p(x)) R3 is the coordinates of p(x) with respect to the basis B for instance

coordV(x+x2) =(0, 1,1)

and

coordB(x+x2)=(0, 0,1)

Also recall the agreement every vector X=(x1, x2, x3) "\isin" R3 is considered as a 3x1 matrix "\begin{pmatrix}
 X1 \\
 X2 \\
 X3 
\end{pmatrix}" i.e, as a column vector

a) Find the matrix MatU"\to" U (T) representing the linear transformation T with respect to the basis of U

b) Verify the equation MatU"\to" U (T)coordU(p(x)) =coordU(T(p(x)), for any p(x) "\isin" R[x] deg≤2

c) Obtain the change of basis matrix PB"\to"U from B to U

d) Hance, or otherwise, obtain coordB(p(x))for each p(x) "\isin" R[x] deg≤2


1
Expert's answer
2021-10-25T18:07:58-0400

a) "T((P(x))=P(x-1)"


"T(1)=0 , T(x)=x-1,"

"T(x)" 2 "=x" 2"-1"

With respect to U

"0= a(1)+b(x)+c(x" 2")"

"a=0\:,b=0,c=0 ,"

"[0]_u =[0,0,0]^T"


"x-1=a(1)+b(x)+c(x^2)"

"a=-1,b=1,c=0"

"[x-1]_u=[-1,1,0]^T"


"x^2-1=a(1)+b(x)+c(x^2)"

"a=-1,b=0,c=1"

"[x^2-1]_u=[-1,0,1]^T"


Matrix "_U" "\to" u(T)"=\begin{pmatrix}
 0&-1 & -1\\
 0&1 & 0\\
0&0&1
\end{pmatrix}"


Part B

Coordu"(P(x))= \begin{bmatrix}
 0 \\
 1 \\
0
\end{bmatrix}"


Coord u"(T(P(x)))="

Coordu "(P(x-1))=\begin{bmatrix}
 -1 \\
 1\\
0
\end{bmatrix}"



"\begin{pmatrix}
 0&-1 & -1\\
 0&1 & 0\\
0&0&1
\end{pmatrix}" "\begin{pmatrix}
 0\\
 1\\
0
\end{pmatrix}" ="\begin{pmatrix}
 -1 \\
 1 \\
0
\end{pmatrix}"


Verified


Part C

Expressing elements of B in terms of the elements of U

"1+x+x^2=a(1)+b(x)+c(x^2)"

"a=1,b=1,c=1"

"[1+x+x^2]_u=[1,1,1]^T"


"2x+x^2=a(1)+b(x)+c(x^2)"

"a=0,b=2,c=1"

"[2x+x^2]_u=[0,2,1]^T"


"x+x^2=a(1)+b(x)+c(x^2)"

"a=0,b=1,c=1"

"[x+x^2]_u=[0,1,1]^T"


Change of base matrix "P_B\to_U="


"\begin{bmatrix}
 1&0 & 0\\
 1&2 & 1\\
1&1&1
\end{bmatrix}"


Part D

Expressing elements of Coordu "(P(x))" in terms of basis B

"x=a(1+x+x^2)+b(2x+x^2)"

"+c(x+x^2)"

"x=a+(a+2b+c)x+(a+b+c)x^2"

"\implies a=0,a+2b+c=1"

"\implies 2b+c=1.......(i)"

"a+b+c=0"

"\implies b+c=0......(ii)"

Solving "(i)" and"(ii)"

"b=1,c=-1"

Coord"_B(P(x))=\begin{pmatrix}
 0 \\
 1\\
-1
\end{pmatrix}"


Alternatively (method 2)


Coordu "(P(x))=\begin{pmatrix}
 0 \\
 1 \\
0
\end{pmatrix}"

Change of basis matrix "P_u\to_B="


"(Matrix P_B\to_u)^{-1}"


"\therefore \begin{pmatrix}
 1&0 & 0 \\
 1&2 & 1\\
1&1&1
\end{pmatrix}" "^-1" ( coord"_B(P(x))")


"\begin{pmatrix}
 1&0& 0\\
 0&1&-1\\
-1&-1&2
\end{pmatrix}" "\begin{pmatrix}
 0 \\
 1 \\
0
\end{pmatrix}" ="\begin{pmatrix}
 0\\
 1\\
-1
\end{pmatrix}"